OpenStudy (anonymous):

The base of a parallelogram is four times its height. If the area of the parallelogram is 324 cm², find the base and height. Enter the numerical answers only with the base first, separated by a comma and a space like this; 1, 2

OpenStudy (anonymous):

So the base is 45, right?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

nevermind Cal.. I already got it my friend showed me.. but Im still confuse tho...

OpenStudy (anonymous):

Ok. The formula for the area of a parallelogram is bh. The base is 4x. The height is x. \[A = 4x \times x \implies 4x^2\]\[342 = 4x^2\]\[81 = x^2\]\[x = 9\]Therefore, 4x = base 4(9) = base = 36 height = x = 9

OpenStudy (anonymous):

yeah I got that too.. well I didnt get that answer, my friend showed me.. lol

OpenStudy (anonymous):

Thanks Cal

OpenStudy (anonymous):

np :)